OpenStudy (anonymous):

if 2(1,2)+x(2,5)=(8,19) what is the value of x

OpenStudy (anonymous):

open up parenthesis. (2,4)+(2x,5x)=(8,19) (2+2x , 4+5x)=(8,19) comparing 2+2x=8 =>2x=8-2 =>2x=6 =>x=3 or 4+5x=19 =>5x=19-4 =>5x=15 =>x=3 so x=3

OpenStudy (anonymous):

thanks!